OpenStudy (anonymous):

square root of 4x-1=1/2

OpenStudy (zpupster):

\[\sqrt{4x-1}=\frac{ 1 }{ 2 }\] lik this?

OpenStudy (zpupster):

hello u there?

OpenStudy (anonymous):

Yes like that,

OpenStudy (zpupster):

ok just solve like other equations doing the same thing to both sides here we will square both sides

OpenStudy (anonymous):

So it would be 4x-1=1/4?

OpenStudy (zpupster):

\[(\sqrt{4x-1})^{2} = (\frac{ 1 }{ 2 })^{2}\]

OpenStudy (zpupster):

leaving us with: 4x-1 = 1/4

OpenStudy (zpupster):

can you solve from here

OpenStudy (anonymous):

I got 5/16?

OpenStudy (zpupster):

yep good job

OpenStudy (anonymous):

Thanks

OpenStudy (zpupster):

always check the work !!